OpenStudy (liizzyliizz):

3√x-2-8=8 (I'm trying to solve for x, but I'm not sure if I'm doing this correctly.)

OpenStudy (liizzyliizz):

( by the way its square root of x-2.. I'm not sure if it shows that clearly.)

jimthompson5910 (jim_thompson5910):

\[\large 3\sqrt{x-2}-8=8\] \[\large 3\sqrt{x-2}=8+8\] \[\large 3\sqrt{x-2}=16\] \[\large \sqrt{x-2}=\frac{16}{3}\] \[\large x-2=(\frac{16}{3})^2\] \[\large x-2=\frac{256}{9}\] \[\large x=\frac{256}{9}+2\] \[\large x=\frac{274}{9}\]

OpenStudy (liizzyliizz):

Thank you! Now that I'm looking at this, I feel silly this is easy -_-